A man invests N1800 at simple interest. After 2 years th amount (including interest) is N1980. Fine the rate perbent per annum
Answer by Theo(13342)

simple interest formula ia:

i = p * r * n

i = interest
p = principal
r = rate of interest per time period
n = number of time periods

man invests 1800
after 2 years he has 1980

his interest earned is 1980 - 1800 = 180.

you get:

p = 1800
i = 180
n = 2


formula of i = p * r * n becomes:

180 = 1800 * r * 2

solve for r to get r = 180 / 1800 / 2 = .05

his interest rate is .05 per year.

that's 5% per year.

at 5% interest per year, he makes 1800 * .05 = 90 interest each year.

in 2 years that becomes 180.

with simple interest, you are only earning interest on the original principal each other.

this gives you the same amount of interest each year.